OpenStudy (anonymous):

(x + y)8, if x = 0.3 and y = 0.7, what is the numerical value of the third term? Possible Answers 0.010 0.020 0.058 0.060 0.068

OpenStudy (welshfella):

(x + y)^8 = x^8 + 8C1 x^7 y + can you continue?

OpenStudy (anonymous):

no i don't understand how to do it at all can you show me the steps

OpenStudy (welshfella):

yes this is the binomial theorem first 3 terms of (x + y)^n = a^n + nC1 x^(n-1) y ^ 1 + nC2 x^(n-2) y^2 so uou are interested in nC2 x^(n-2) y^2 so plug in the values of n, x and y and you have the answer